# Metric theory of inhomogeneous Diophantine approximations with a fixed matrix

### v2

@@ 6.2 Statements about measure
The following lemmata will be used in the proofs of our main statements.
@@
There exists an infinitely small on infinity function $\xi(t)$ such that for any $\frac{1}{2}>\sigma>0$ one has
$\frac{m\left(U(\sigma)\cap\left(B_{M}(0)+{\bf\mathfrak{a}}\right)\right)}{m(B_{M}(0))}=2\sigma\cdot(1+\xi({M}\cdot{|pr_{\mathcal{L}}{\bf u}|})),$
where $B_{M}(0)$ denotes the ball ( with respect to sup-norm in $\mathbb{R}^{d}$) of radius $M$ intersected with $\mathcal{L}$, $m$ is $(\dim\mathcal{L})$-dimensional Lebesgue measure, and ${\bf\mathfrak{a}}\in\mathcal{A}$ is any vector.
@@
Let ${\bf v}_{k}\in\mathcal{L}^{\perp}$ be defined by ${\bf u}=pr_{\mathcal{L}}{\bf u}+{\bf v}_{k}$. Let $\xi_{k}=\{{\bf\mathfrak{a}}\cdot{\bf u}\}\in[0,1)$. Then, for any $\boldsymbol{\eta}=\tilde{\boldsymbol{\eta}}+{\bf\mathfrak{a}}\in\mathcal{A}$,
@@
where the shift vector $\xi_{k}\frac{pr_{\mathcal{L}}{\bf u}}{|pr_{\mathcal{L}}{\bf u}|^{2}}$ has its length bounded from above by $\frac{1}{\mathfrak{D}}$. By this,
@@
Fix $M>0$. Lemma 4 shows that the bigger $|pr_{\mathcal{L}}{\bf u}|$ gets, the closer $U$ and any ball of radius $M$ are to independent sets. This provides us with the following useful claim:
@@
$\lim\limits_{k,s\rightarrow\infty,\,\,k\neq s}\frac{m\left(U_{k}\cap U_{s}\cap\left(B_{M}(0)+{\bf\mathfrak{a}}\right)\right)}{m\left(U_{s}\cap\left(B_{M}(0)+{\bf\mathfrak{a}}\right)\right)m\left(U_{k}\cap\left(B_{M}(0)+{\bf\mathfrak{a}}\right)\right)}=m\left(B_{M}(0)\right).$
@@
Let $\omega(t)$ be such a continuously decreasing function that $\sum\limits_{\nu=1}^{\infty}\omega(Y_{\nu})<\infty$. Let $\mathcal{A}$ be an non-1-exceptional affine subspace of $\mathbb{R}^{n}$ of positive dimension. For almost all $\boldsymbol{\eta}\in\mathcal{A}$, one has
@@
The property of being not $1$-exceptional guarantees that there exists such a $\mathfrak{D}>0$ that $|\widehat{{\bf y}_{\nu},\mathcal{L}}-\frac{\pi}{2}|>\frac{\mathfrak{D}}{|{\bf y}_{\nu}|}$ for all $\nu$ (except maybe finitely many). This guarantees that $|pr_{\mathcal{L}}{\bf y}_{\nu}|>\mathfrak{D}$. By Lemma 4,
@@
Another classical result which will be used in the proofs is the following lemma:
@@
(Lemma 5, §3, Ch.1 in [47]) Let $X$ be a measure space with measure $\mu$, and let $\{E_{k}\}_{k=1}^{\infty}$ be a countable collection of measurable sets. Let $E=\bigcap\limits_{N=1}^{\infty}\bigcup\limits_{k=N}^{\infty}E_{k}$. Suppose $\sum\limits_{k=1}^{\infty}\mu(E_{k})=\infty$. Then,
@@ 7.1 Some general theorems via transference
In this section, we will formulate and prove general results which imply Theorems A, 2, B, 4 and 6. The deduction is provided by setting
$f(T)=\psi^{m}\left(T^{\frac{1}{n}}\right)\,\,\,\,\,\,\,\,\text{and}\,\,\,\,\,\,\,\,\,\,\,\,g(T)=\frac{1}{\rho^{n}\left(T^{\frac{1}{m}}\right)}.$ (26)
@@
$\rho(t)\,\,\,\text{be the function inverse to the function}\,\,\,t\mapsto 1/\psi(t).$ (29)
@@
Theorem 1 immediately follows from Lemma I.
Proof.
of Theorem 1. Let $Y$ be such a number that
$\psi_{\Theta^{\top}}(Y)\geq\psi(Y).$ (30)
Let us define $Q$ by $\frac{\kappa}{Q}=\psi(Y)$; then $Y=\rho(\frac{1}{\kappa}Q)$ and by Lemma I there exists ${\bf q}\in\mathbb{Z}^{m}$ with $|{\bf q}|\leq Q$ that
@@
  • (a)
    The set of $Y$ such that (30) holds is unbounded, therefore such is the set of $Q$ for which (31) holds, which completes the proof.
  • (b)
    Condition (30) holds for any $Y$ large enough, so (31) holds for all $Q$ large enough.
@@
    The inequality
    ${\psi_{\Theta^{\top}}(t)}\leq{\psi(t)}$ (32)
    holds for all $t\in\mathbb{R}_{+}$ large enough, or
@@
    the inequality
    ${\psi_{\Theta^{\top}}(t)}\leq{\psi(t)}$ (33)
    holds for an unbounded set of $t\in\mathbb{R}_{+}$.
@@
Let $\Theta^{\top}$ be not trivially singular, and let $\psi$ be a continuous function decreasing to zero as $t\rightarrow+\infty$, and such that the inequality (32) holds for all $t\in\mathbb{R}_{+}$ large enough.
@@
Let $\omega(t)$ be such a function that $\sum\limits_{\nu=1}^{\infty}\omega(Y_{\nu})<\infty$. Let $G(t)$ be the inverse function for $\frac{\omega(t)}{\psi(t)}$ and let $\tilde{\rho}(t):=\frac{G(t)}{\omega\left(G(t)\right)}$. Then, for almost any $\boldsymbol{\eta}\in\mathcal{A}$ the inequality
@@
be some family of non-increasing functions, parametrized by $\boldsymbol{\eta}\in\mathbb{R}^{n}$. We also want $F_{\boldsymbol{\eta}}(t)$ to decrease slower than $\psi(t)$, so that the function $\frac{F_{\boldsymbol{\eta}}(t)}{\psi(t)}$ is increasing to infinity. Let $G_{\boldsymbol{\eta}}(t)$ be the inverse function for $\frac{F_{\boldsymbol{\eta}}(t)}{\psi(t)}$. Let $\tilde{\rho}_{\boldsymbol{\eta}}(t):=\frac{G_{\boldsymbol{\eta}}(t)}{F_{\boldsymbol{\eta}}\left(G_{\boldsymbol{\eta}}(t)\right)}$.
@@
Let us also note that if $\Theta^{\top}$ is trivially singular and $\Theta^{\top}{\bf z}={\bf 0}$ for some ${\bf 0\neq z}\in\mathbb{Z}^{n}$, (34) still holds with ${\bf y}_{\nu}$ replaced by $\bf z$}. Thus,
@@
The next lemma is based on this simple argument.
@@
Suppose $\Theta$ and $\psi$ are chosen in such a way that (32) or (33) holds. Suppose $\boldsymbol{\eta}\in\mathbb{R}^{n}$ satisfies the condition $\boldsymbol{\eta}\cdot{\bf y}_{\nu}||\geq F_{\boldsymbol{\eta}}(Y_{\nu})$ for any $\nu$ large enough. Then the condition
@@
    Let us fix ${\bf q}\in\mathbb{Z}^{m}$ and choose such $\nu$ that
    $\frac{F_{\boldsymbol{\eta}}(Y_{\nu})}{\psi(Y_{\nu})}\leq 2m|{\bf q}|<\frac{F_{\boldsymbol{\eta}}(Y_{\nu+1})}{\psi(Y_{\nu+1})};$ (36)
@@
    $2m|{\bf q}|<\frac{F_{\boldsymbol{\eta}}(Y_{\nu+1})}{\psi(Y_{\nu+1})}\leq\frac{F_{\boldsymbol{\eta}}(Y_{\nu})}{\psi(Y_{\nu+1})}.$ (37)
    Let us notice that by (32) for any $\nu$ large enough one has $\psi_{\Theta^{\top}}(Y_{\nu})\leq\psi(Y_{\nu+1})$.
    Using this inequality, (34) and (37), we get that
@@
    $\Theta{\bf q}-\boldsymbol{\eta}||\geq\frac{1}{2n}\frac{F_{\boldsymbol{\eta}}(Y_{\nu})}{Y_{\nu}}.$ (38)
    Applying $G_{\boldsymbol{\eta}}$ to the first inequality in (36), one can see that
@@
    By (33), there exists infinitely many $\nu$ such that
    $\psi_{\Theta^{\top}}(Y_{\nu})\leq\psi(Y_{\nu}).$ (39)
@@
    $2m|{\bf q}|\leq\frac{F_{\boldsymbol{\eta}}(Y_{\nu})}{\psi(Y_{\nu})}:=2mQ_{\nu}.$ (40)
    Using (34), (40) and (39), we get that
@@
    which again implies (38).
    We apply $G_{\boldsymbol{\eta}}$ to (40) to see that
@@
To prove Theorem 2, we will need the following statement, which is a special case of the Theorem 4.1 in [10].
@@
Let $\{{\bf z}_{\nu}\}_{\nu\in\mathbb{N}}$ be a sequence of real $d$-vectors with their norms forming a lacunary sequence. Then the set
@@
Let $\widehat{{\bf v},V}$ denote the angle between the vector $\bf v\in\mathbb{R}^{n}$ and linear space $V$. The set $\Omega$ is compact, so there exists
$0<\varphi_{0}=\inf\limits_{v\in\Omega}|\widehat{{\bf v},\mathcal{L}}-\frac{\pi}{2}|=\min\limits_{v\in\Omega}|\widehat{{\bf v},\mathcal{L}}-\frac{\pi}{2}|.$
Let $\Phi_{k}:=\{v\in\mathbb{R}^{n}:\,\,\,\left(\frac{2}{3}\right)^{k}<\cos\left(\widehat{v,\mathcal{L}}\right)\leq\left(\frac{2}{3}\right)^{k-1}\}$. Then, there exists $k_{0}$ such that $\Omega\subset\bigcup\limits_{k=1}^{k_{0}}\Phi_{k}$, or, in other words, for any $\nu$ there exists $k\in\{1,\ldots,k_{0}\}$ such that ${\bf y}_{\nu}\in\Phi_{k}$. Now, let us represent the sequence $\{{\bf y}_{\nu}\}$ as the union of finitely many subsequences as follows:
  • Let ${\bf y}_{\nu}^{j}:={\bf y}_{A(\nu-1)+j},\,\,\,j=1,\ldots,A,$ where $A$ is as in Proposition 1. Then
    $\{{\bf y}_{\nu}\}=\bigcup\limits_{j=1}^{A}\{{\bf y}_{\nu}^{j}\}\,\,\,\,\,\,\,\,\,\text{and}\,\,\,\,\,\,\,\,\,\frac{|{\bf y}_{\nu+1}^{j}|}{|{\bf y}_{\nu}^{j}|}\geq 2\,\,\,\,\,\text{ for any}\,\,\nu.$
@@
    $\frac{\Big{|}\left({\bf y}_{\nu+1}^{j,k}\right)^{\mathcal{L}}\Big{|}}{\Big{|}\left({\bf y}_{\nu}^{j,k}\right)^{\mathcal{L}}\Big{|}}=\frac{\Big{|}{\bf y}_{\nu+1}^{j,k}\Big{|}\cos\left(\widehat{|{\bf y}_{\nu+1}^{j,k},\mathcal{L}}\right)}{\Big{|}{\bf y}_{\nu}^{j,k}\Big{|}\cos\left(\widehat{|{\bf y}_{\nu}^{j,k},\mathcal{L}}\right)}\geq\frac{2}{3}\frac{\Big{|}{\bf y}_{\nu+1}^{j,k}\Big{|}}{\Big{|}{\bf y}_{\nu}^{j,k}\Big{|}}\geq\frac{4}{3};$ (41)
    in particular, each of the sequences $\{({\bf y}_{\nu}^{j,k})^{\mathcal{L}}\}$ is lacunary.
@@
    It’s easy to see that $N=\bigcap\limits_{j=1}^{A}\bigcap\limits_{k=1}^{k_{0}}N(j,k)$, so it’s enough to prove that each of the sets $N(j,k)\cap\mathcal{A}$ is winning. By (41), this follows from Proposition 6.

### v3

@@ 6.2 Statements about measure
The following lemmata will be used in proofs of our main statements. We denote the orthogonal projection of a vector $\bf u$} on a linear subspace $\mathcal{L}$ of $\mathbb{R}^{n}$ by $pr_{\mathcal{L}}{\bf u}$.
@@
There exists a function $\xi(t)$ satisfying $\lim\limits_{t\rightarrow\infty}\xi(t)=0$ such that for any $\frac{1}{2}>\sigma>0$ one has
$\frac{m\left(U(\sigma)\cap\left(B_{M}(0)+{\bf{\bf a}}\right)\right)}{m(B_{M}(0))}=2\sigma\cdot(1+\xi({M}\cdot{|pr_{\mathcal{L}}{\bf u}|})),$
where $B_{M}(0)$ denotes the ball (with respect to the sup-norm in $\mathbb{R}^{d}$) of radius $M$ intersected with $\mathcal{L}$, $m$ is $(\dim\mathcal{L})$-dimensional Lebesgue measure, and ${\bf{\bf a}}\in\mathcal{A}$ is any vector.
@@
Let $\bf v$$\in\mathcal{L}^{\perp}$ be defined by ${\bf u}=pr_{\mathcal{L}}{\bf u}+{\bf v}$. Let $\xi=\{{\bf{\bf a}}\cdot{\bf u}\}\in[0,1)$ be the fractional part of scalar product of $\bf a$} and $\bf u$}. Then, for any $\boldsymbol{\eta}=\tilde{\boldsymbol{\eta}}+{\bf{\bf a}}\in\mathcal{A}$,
@@
Thus,
@@
Fix $M>0$. Lemma 4 shows that the bigger $|pr_{\mathcal{L}}{\bf u}|$ gets, the closer $U(\sigma)$ and any ball of radius $M$ are to independent sets. This provides us with the following useful claim:
@@
Let $\mathcal{A}$ be an affine subspace of $\mathbb{R}^{n}$ with $\mathcal{L}$ as the corresponding linear subspace, and let the sequence $\{{\bf u}_{k}\}$ of vectors in $\mathbb{R}^{n}$ be such that $\lim\limits_{k\rightarrow\infty}\frac{|pr_{\mathcal{L}}{\bf u}_{k}|}{|pr_{\mathcal{L}}{\bf u}_{k+1}|}=0$. Let
@@
Fix $\sigma<\frac{1}{2}$. Then, for any $M>0$ and any ${\bf a}\in\mathcal{A}$, one has
$\lim\limits_{k,s\rightarrow\infty,\,\,k\neq s}\frac{m\left(U_{k}\cap U_{s}\cap\left(B_{M}(0)+{\bf{\bf a}}\right)\right)}{m\left(U_{s}\cap\left(B_{M}(0)+{\bf{\bf a}}\right)\right)m\left(U_{k}\cap\left(B_{M}(0)+{\bf{\bf a}}\right)\right)}=m\left(B_{M}(0)\right).$
@@
Let $\omega(t)$ be a continuous decreasing function such that $\sum\limits_{\nu=1}^{\infty}\omega(Y_{\nu})<\infty$. Let $\mathcal{A}$ be an non-1-exceptional affine subspace of $\mathbb{R}^{n}$ of positive dimension. For almost all $\boldsymbol{\eta}\in\mathcal{A}$, one has
@@
The property of being not $1$-exceptional guarantees that there exists such a $\delta>0$ that $|\measuredangle({{\bf y}_{\nu},\mathcal{L}})-\frac{\pi}{2}|>\frac{\delta}{|{\bf y}_{\nu}|}$ for all $\nu$ (except maybe finitely many). This guarantees that $|pr_{\mathcal{L}}{\bf y}_{\nu}|>\delta$. By Lemma 4,
@@
Another classical result which will be used in the proofs is Chung-Erdös inequality (see [12]; we are using it in form of Lemma 5, §3, Ch.1 in [49]).
@@
Let $(X,\mu)$ be a measure space, and let $\{E_{k}\}_{k=1}^{\infty}$ be a countable collection of measurable sets. Let $E=\bigcap\limits_{N=1}^{\infty}\bigcup\limits_{k=N}^{\infty}E_{k}$. Suppose $\sum\limits_{k=1}^{\infty}\mu(E_{k})=\infty$. Then,
@@ 7.1 Some general theorems via transference
In this section we will formulate and prove general results which imply Theorems A, 2, B, 4 and 6. The deduction is provided by setting
$f(T)=\psi^{m}\left(T^{\frac{1}{n}}\right)\,\,\,\,\,\,\,\,\text{and}\,\,\,\,\,\,\,\,\,\,\,\,g(T)=\frac{1}{\rho^{n}\left(T^{\frac{1}{m}}\right)}.$ (7.1)
@@
$\rho(t)\,\,\,\text{be the function inverse to the function}\,\,\,t\mapsto 1/\psi(t).$ (7.4)
@@
Theorem 1 is a consequence of Lemma I.
Proof of Theorem 1. Let $Y>0$ be such that
$\psi_{\Theta^{\top}}(Y)\geq\psi(Y).$ (7.5)
Let us define $t$ by $\frac{\kappa}{t}=\psi(Y)$; then $Y=\rho(\frac{1}{\kappa}t)$ and, by Lemma I, there exists ${\bf q}\in\mathbb{Z}^{m}$ with $|{\bf q}|\leq t$ such that
@@
  1. (a)
    The set of $Y$ such that (7.5) holds is unbounded, therefore so is the set of $t$ for which (7.6) holds, which completes the proof.
  2. (b)
    Condition (7.5) holds for any $Y$ large enough, so (7.6) holds for all $t$ large enough, and thus we get the desired statement.
@@
    $\text{the inequality}\,\,{\psi_{\Theta^{\top}}(t)}\leq{\psi(t)}\,\,\text{holds for all $t\in\mathbb{R}_{+}$ large enough,}$ (7.7)
    or
@@
    $\text{the inequality}\,\,{\psi_{\Theta^{\top}}(t)}\leq{\psi(t)}\,\,\text{holds for an unbounded set of $t\in\mathbb{R}_{+}$.}$ (7.8)
@@
Let $\mathcal{A}$ be an affine subspace of $\mathbb{R}^{n}$ of positive dimension, with $\mathcal{L}$ as the corresponding linear subspace and $\mathcal{L}^{\perp}$ its orthogonal complement. If $\Theta^{\top}$ is trivially singular, assume also that $\mathcal{A}$ is not exceptional.
@@
Let $\Theta^{\top}$ be not trivially singular, and let $\psi$ be a continuous function decreasing to zero as $t\rightarrow+\infty$, such that (7.7) holds for all $t\in\mathbb{R}_{+}$ large enough.
@@
Let $\omega(t)$ be a function such that $\sum\limits_{\nu=1}^{\infty}\omega(Y_{\nu})<\infty$. Let $G(t)$ be the inverse function to $\frac{\omega(t)}{\psi(t)}$, and let $\tilde{\rho}(t):=\frac{G(t)}{\omega\left(G(t)\right)}$. Then, for almost any $\boldsymbol{\eta}\in\mathcal{A}$ the inequality
@@
Remark 0.
We note that Theorem 6 follows from Theorem 3 via the identifications
@@
be some family of non-increasing functions, parametrized by $\boldsymbol{\eta}\in\mathbb{R}^{n}$. We also want $F_{\boldsymbol{\eta}}(t)$ to decrease slower than $\psi(t)$, so that the function $\frac{F_{\boldsymbol{\eta}}(t)}{\psi(t)}$ is increasing to infinity. Let $G_{\boldsymbol{\eta}}(t)$ be the inverse function to $\frac{F_{\boldsymbol{\eta}}(t)}{\psi(t)}$. Let $\tilde{\rho}_{\boldsymbol{\eta}}(t):=\frac{G_{\boldsymbol{\eta}}(t)}{F_{\boldsymbol{\eta}}\left(G_{\boldsymbol{\eta}}(t)\right)}$.
@@
Let us also note that if $\Theta^{\top}$ is trivially singular and $\Theta^{\top}{\bf z}={\bf 0}$ for some ${\bf 0\neq z}\in\mathbb{Z}^{n}$, (7.9) still holds with ${\bf y}_{\nu}$ replaced by $\bf z$}. Thus,
@@
Suppose $\Theta$ and $\psi$ are chosen in such a way that
  • (a)
    (7.7) holds or
  • (b)
    (7.8) holds.
Suppose $\boldsymbol{\eta}\in\mathbb{R}^{n}$ satisfies the condition $\boldsymbol{\eta}\cdot{\bf y}_{\nu}||\geq F_{\boldsymbol{\eta}}(Y_{\nu})$ for any $\nu$ large enough. Then the condition
@@
    Let us fix ${\bf q}\in\mathbb{Z}^{m}$ with large enough norm, such that $\frac{F_{\boldsymbol{\eta}}(Y_{1})}{\psi(Y_{1})}\leq 2m|{\bf q}|$. Choose $\nu$ such that
    $\frac{F_{\boldsymbol{\eta}}(Y_{\nu})}{\psi(Y_{\nu})}\leq 2m|{\bf q}|<\frac{F_{\boldsymbol{\eta}}(Y_{\nu+1})}{\psi(Y_{\nu+1})};$ (7.11)
@@
    $2m|{\bf q}|<\frac{F_{\boldsymbol{\eta}}(Y_{\nu+1})}{\psi(Y_{\nu+1})}\leq\frac{F_{\boldsymbol{\eta}}(Y_{\nu})}{\psi(Y_{\nu+1})}.$ (7.12)
    Let us notice that by (7.7) for any $\nu$ large enough one has $\psi_{\Theta^{\top}}(Y_{\nu})\leq\psi(Y_{\nu+1})$.
    Using this inequality, (7.9) and (7.12), we get that
@@
    $\Theta{\bf q}-\boldsymbol{\eta}||\geq\frac{1}{2n}\frac{F_{\boldsymbol{\eta}}(Y_{\nu})}{Y_{\nu}}.$ (7.13)
    Applying $G_{\boldsymbol{\eta}}$ to the first inequality in (7.11), one can see that
@@
    By (7.8), there exists infinitely many $\nu$ such that
    $\psi_{\Theta^{\top}}(Y_{\nu})\leq\psi(Y_{\nu}).$ (7.14)
@@
    $2m|{\bf q}|\leq\frac{F_{\boldsymbol{\eta}}(Y_{\nu})}{\psi(Y_{\nu})}:=2mQ_{\nu}.$ (7.15)
    Using (7.9), (7.15) and (7.14), we get that
@@
    which again implies (7.13).
    We apply $G_{\boldsymbol{\eta}}$ to (7.15) to see that
@@
To prove Theorem 2, we will need the following statement, which is a special case of the Theorem 4.1 in [10]. We denote Euclidean norm by $|\cdot|_{2}$.
@@
Let $\{{\bf z}_{\nu}\}_{\nu\in\mathbb{N}}$ be a sequence of vectors in $\mathbb{R}^{d}$ such that the sequence $\{|{\bf z}_{\nu}|_{2}\}_{\nu\in\mathbb{N}}$ of their Euclidean norms is lacunary. Let $\{{\delta}_{\nu}\}_{\nu\in\mathbb{N}}$ be a sequence of real numbers. Then the set
@@
Let $\measuredangle({{\bf v},\mathcal{L}})$ denote the angle between the vector $\bf v\in\mathbb{R}^{n}$ and linear subspace $\mathcal{L}$. The set $\Omega$ is compact, so there exists
$0<\varphi_{0}=\inf\limits_{v\in\Omega}\left|\measuredangle({{\bf v},\mathcal{L}})-\frac{\pi}{2}\right|=\min\limits_{v\in\Omega}\left|\measuredangle({{\bf v},\mathcal{L}})-\frac{\pi}{2}\right|.$
Let $\Phi_{k}:=\{v\in\mathbb{R}^{n}:\,\,\,\left(\frac{2}{3}\right)^{k}<\cos\left(\measuredangle({v,\mathcal{L}})\right)\leq\left(\frac{2}{3}\right)^{k-1}\}$. Then, there exists $k_{0}$ such that $\Omega\subset\bigcup\limits_{k=1}^{k_{0}}\Phi_{k}$, or, in other words, for any $\nu$ there exists $k\in\{1,\ldots,k_{0}\}$ such that ${\bf y}_{\nu}\in\Phi_{k}$. Now, let us represent the sequence $\{{\bf y}_{\nu}\}$ as the union of finitely many subsequences as follows:
  • Let ${\bf y}_{\nu}^{j}:={\bf y}_{C(\nu-1)+j},\,\,\,j=1,\ldots,C,$ where $C$ is such a constant that $\frac{|{\bf y}_{\nu+C}|_{2}}{|{\bf y}_{\nu}|_{2}}>2$; the existence of such $C$ follows from Proposition 1, since sup norm is equivalent to Euclidean norm. Then
    $\{{\bf y}_{\nu}\}=\bigcup\limits_{j=1}^{C}\{{\bf y}_{\nu}^{j}\}\,\,\,\,\,\,\,\,\,\text{and}\,\,\,\,\,\,\,\,\,\frac{|{\bf y}_{\nu+1}^{j}|_{2}}{|{\bf y}_{\nu}^{j}|_{2}}\geq 2\,\,\,\,\,\text{ for any}\,\,\nu.$
@@
    $\frac{\Big|pr_{\mathcal{L}}{\bf y}_{\nu+1}^{j,k}\Big|_{2}}{\Big|pr_{\mathcal{L}}{\bf y}_{\nu}^{j,k}\Big|_{2}}=\frac{\Big|{\bf y}_{\nu+1}^{j,k}\Big|_{2}\cos\left(\measuredangle({{\bf y}_{\nu+1}^{j,k},\mathcal{L}})\right)}{\Big|{\bf y}_{\nu}^{j,k}\Big|_{2}\cos\left(\measuredangle({{\bf y}_{\nu}^{j,k},\mathcal{L}})\right)}\geq\frac{2}{3}\frac{\Big|{\bf y}_{\nu+1}^{j,k}\Big|_{2}}{\Big|{\bf y}_{\nu}^{j,k}\Big|_{2}}\geq\frac{4}{3};$ (7.16)
    in particular, each of the sequences $\{|pr_{\mathcal{L}}{\bf y}_{\nu}^{j,k}|_{2}\}$ is lacunary.
@@
    It is easy to see that $N=\bigcap\limits_{j=1}^{A}\bigcap\limits_{k=1}^{k_{0}}N(j,k)$, so it is enough to prove that each of the sets $N(j,k)\cap\mathcal{A}$ is HAW. This follows from (7.16) and Proposition 7 by a standard argument:
  • Fix indices $j$ and $k$. Fix a vector ${\bf a}\in\mathcal{A}$, and let $\delta_{\nu}:={\bf a}\cdot{\bf y}_{\nu}^{j,k}$. Let ${\bf l}_{1},\ldots,{\bf l}_{s}$ be an orthonormal basis of $\mathcal{L}$, and $L$ a matrix with columns ${\bf l}_{i},\,\,i=1,\ldots,s$. Then, $\boldsymbol{\lambda}\mapsto L\boldsymbol{\lambda}+{\bf a}$ is an affine bijection between $\mathbb{R}^{s}$ and $\mathcal{A}$. Thus, it is enough to show that the set
[…]
    is HAW.
    Since the sequence $\{|pr_{\mathcal{L}}{\bf y}_{\nu}^{j,k}|_{2}\}$ is lacunary, so is the sequence $\left\{\left|\left({\bf y}_{\nu}^{j,k}\right)^{\top}L\right|_{2}\right\}_{\nu\in\mathbb{N}}$. Now HAW property of the set (7.17) follows directly from Proposition 7.